DIRC Operations Report

Wednesday 2008/01/09 to Wednesday 2008/01/16


Oncall: Jose Benitez (Joe Schwiening and Jennifer Prendki on backup)
Fast Monitoring: Joe Schwiening, Jose Benitez, Jennifer Prendki

Contents:

    PEP Activities     BaBar Activities     DIRC activities     Backgrounds     Scheduled activities     Pending Work     Links

Luminosities this week

PEP peak lumi: ~9.9 x 1033 (last week: 8.3 x 1033)

PEP delivery: 2.2 fb-1

BaBar integrated lumi:2.9 fb-1

Run 7 total delivery: 10.55 fb-1

Run 7 total BaBar recorded: 10.44 fb-1 total, including 0 fb-1 off-peak data and including 9.82 fb-1 on the Y(3S)

PEP activities/issues this week

  • There was a ROD on Wednesday and Thursday because PEP needed to replace a kicker magnet on the HER.
  • "BOAT_LER" system broke in the swing shift. Caused some down time.

    BaBar activities/issues this week

  • Many subsystems used the ROD on Wednesday and Thursday to make minor repairs.
  • Usually smooth data taking.

    DIRC activities/issues this week

  • During the ROD on Jan 9: There was a new flow sensor installed on the chiller. This appears to be working fine, no glitches.
  • Also, Matt,Jerry and Jose made a controlled access to the top of the detector to investigate the flaky Nitrogen flow meter electronics. No problem was found. The electronics were just cleaned. The instabilities have not returned.
  • One channel was masked because it became noisy. Sector 0, DFB 5, channel 2.
  • On Thursday evening there was an automatic page due to the circulation pump on the water plant. This was found to be due to a glitch on the building power.
  • Friday we started to see bad timing distributions on the Fast Mon Plots. There were 3 periods where the distributions were so bad that the signal peak was not visible. This bad distributions appeared even though the scalers and other subsystems were OK on backgrounds. This has now been found to be due to bad trickle injection on the LER. This problem has persisted.
  • One HV channel in Sector 2 tripped and was ramped back up shortly after. The trip appears to be due to bad backgrounds, the channel looks healthy.

    See DIRC logbook for details

    Lumi and Background this week